OpenStudy (anonymous):

by using the remainder theorem is 2 a zero of 4x^3-5x^2+3x-10?

OpenStudy (amistre64):

the remainder thrm is quite .... umm .... self evident i think if the remainder is zero; then the proposed value is a zero of the function

OpenStudy (amistre64):

do we synth it or just plug it in?

OpenStudy (amistre64):

4*2 = 8-5=3*2=6+3=9*2=18-10=8 not 0

OpenStudy (anonymous):

i wanted to use synthetic division but i couldnt figure out what number i would use to multiply everything by

OpenStudy (amistre64):

2 they tell you "2" :)

OpenStudy (anonymous):

lol oh wow, so its a -2 sincei have to take the opposite sign?

OpenStudy (amistre64):

nah, you want to know if 2 is a zero; or rather, is (x+2) a factor

OpenStudy (amistre64):

(x-2) typoed it lol

OpenStudy (anonymous):

so then how should i set up my problem?

OpenStudy (amistre64):

pull of the coeffs; and since we are all accounted for there is no fill in 0s to be had

OpenStudy (amistre64):

4x^3 -5x^2 +3x -10 3 2 1 0 4 -5 3 -10 0 <--- add up ------------ 2) 4 <--- multiply

OpenStudy (amistre64):

4*2=8-5=3*2=6+3=9*2=18-10=8 not 0

OpenStudy (anonymous):